OpenStudy (vera_ewing):

How many moles of magnesium chloride are formed when 55 mL of 0.70 M HCl is added to 42 mL of 1.204 g Mg(OH)2? Mg(OH)2(s) + 2HCl(aq) MgCl2(aq) + 2H2O(l) A. 0.021 mol B. 0.019 mol C. 0.012 mol D. 0.039 mol

OpenStudy (vera_ewing):

@aaronq Can you please help me with this question?

OpenStudy (aaronq):

are you sure you copied the question right? this "42 mL of 1.204 g Mg(OH)" doesnt make sense

OpenStudy (vera_ewing):

Well it's supposed to be Mg(OH)2

OpenStudy (aaronq):

i mean this part "42 mL of 1.204 g"

OpenStudy (vera_ewing):

Yep it's right...

OpenStudy (aaronq):

do you think they meant "42 mL of 1.204 M" or "added to 1.204 g Mg(OH)2" you can't add 42 mL of a solid

OpenStudy (vera_ewing):

yeah I think that's what they meant.

OpenStudy (aaronq):

the second right? so you would first find the limiting reagent. to do this you convert both reactants to moles, then divide each by it's coefficient in the balanced equation. Whichever has less after this division is the limting reagent

OpenStudy (vera_ewing):

Ok so the answer is C right?

OpenStudy (aaronq):

idk, i havent worked it out. i can check your steps if you post them

OpenStudy (vera_ewing):

0.038 x (1 mol MgCl2/2 mols HCl) = approx 0.038 x 1/2 = approx 0.019 0.05 x (1 mol MgCl2/1 mol Mg(OH)2) = 0.0 x 1/1 = approx 0.05

OpenStudy (aaronq):

how did you get 0.05?

OpenStudy (vera_ewing):

mols HCl = M x L = approx 0.038 mols Mg(OH)2 = approx 0.05

OpenStudy (aaronq):

the moles of magnesiums hydroxide are 0.0206448 moles

OpenStudy (vera_ewing):

oh oops :/

OpenStudy (aaronq):

so you'd use the one with less moles.. which is HCl

OpenStudy (vera_ewing):

then what?

OpenStudy (aaronq):

then find the moles MgCl2 as you would normally \(\sf \dfrac{moles~of~HCl}{HCl's ~coefficient}=\dfrac{moles~of~MgCl_2}{MgCl_2's ~coefficient}\)

OpenStudy (aaronq):

plug in what you know then solve (for moles of MgCl2)

OpenStudy (vera_ewing):

oh ok so it is C lol. thanks for your help :)

OpenStudy (aaronq):

nope, it's B